Use the References to access important values if neededfor this question.The radioactive isotope strontium-81...

The radioactive isotope strontium-81 is used in medical imaging as indicated on the table above. If 57.7 milligrams of strontium-81 is administered to a patient, how many milligrams are left in the body after 44.4 minutes?

mg
Answer #1

Half life period of strontium \(-81, t_{1}=22.3 \mathrm{~min}\) No of halflives \(44.4\) minutes, \(n=\frac{44.4 \mathrm{~min}}{22.3 \mathrm{~min}}=1.99\)

Initial amount of the strontium \(-81, a=57.7 \mathrm{mg}\) Amount of the strontium \(-81\) left after \(1.99\) halflives \(=\left(\frac{1}{2}\right)^{n} \times \mathrm{a}\)

$$ \begin{aligned} &=\left(\frac{1}{2}\right)^{1.99} \times 57.7 \mathrm{mg} \\ &=(0.252) \times 57.7 \mathrm{mg} \\ &=14.5 \mathrm{mg} \end{aligned} $$